OpenStudy (anonymous):

Which of the following equations describe (1,6) and (-3,-6)

OpenStudy (anonymous):

to find an equation given two points first need to find the slope, m, by using the formula \(\Large m=\frac{y_2-y_1}{x_2-x_1}\) then substitute it in the slope-intercept equation of the line \(y=mx+b\) where b is the y-int. to solve for "b", plug one of the points in the equation, \(y=mx+b\), with the slope, m. Welcome to Openstudy! ☺